# Binary-Report Peer Prediction for Real-Valued Signal Spaces

### v1

@@ 4 Determinant-based Mutual Information (DMI) Mechanism
4.1 Equilibrium Characterization
Unlike DG, where we could compute the payment for a single task in isolation (albeit requiring multiple tasks from a different agent), DMI is inherently a multi-task mechanism. As in the original analysis of DMI, here we restrict to consistent strategies, where the same $\sigma$ is applied to each $X_{i}$. (See below for behavior beyond this case.) Even under this restriction, threshold equilibria are much more complex in DMI given the interactions between four different signals, making it unclear how to reason directly from the definition of a threshold equilibrium. Instead, we restrict all agents to playing threshold strategies. Let
$U_{i}(\sigma,\sigma^{\prime})=\mathop{\mathbb{E}}\left[M_{\textrm{DMI}}(\sigma(X_{1}),\ldots,\sigma(X_{4}),\sigma^{\prime}(X_{1}^{\prime}),\ldots,\sigma^{\prime}(X_{4}^{\prime}))\right].$ (14)
be the (ex-ante) expected utility for playing threshold strategy $\sigma_{i}$ given the other agent using $\sigma_{j}$.
@@ 4.1 Equilibrium Characterization
A threshold strategy $\sigma:\mathbb{R}\to\mathcal{R}$ is a equilibrium restricted to threshold strategies under DMI if for all threshold strategies $\hat{\sigma}:\mathbb{R}\to\mathcal{R}$, $U_{i}(\sigma,\sigma)\geq U_{i}(\hat{\sigma},\sigma)$.
This is a weaker condition than being a threshold equilibrium becacuse it rules out deviations to non-threshold strategies, an issue we will return to. However, by restricting to this space we can provide an equilibrium characterization in the same spirit as we did for OA and DG.
This unbiasedness is implied by Kong [2024, Claim 4.4], which we paraphrase for this binary-report context: for any $Y,Y^{\prime}$, there exists a constant $\alpha>0$ such that $\mathop{\mathbb{E}}\det\sum_{i\in S}\mathbf{1}_{Y}\mathbf{1}_{Y^{\prime}}^{T}=\alpha\det U_{Y,Y^{\prime}}$.
Equilibrium results.
Proposition 5.
$\tau^{*}=\pm\infty$ are both always equilibria restricted to threshold strategies under DG.
Theorem 5.
Proof.
00 0 $=\frac{d}{d\tau}\alpha^{\prime}h(\tau,\tau^{*})^{2}=2h(\tau,\tau^{*})\alpha^{\prime}f(\tau)\left(\Pr[X^{\prime}>\tau^{*}]-\Pr[X^{\prime}>\tau^{*}\mid X=\tau]\right)~{}$ (16)
Theorem 5 shows that DMI has the same necessary condition for equilibrium as DG. So, modulo the slightly different sufficient condition for equilibrium, its equilibria can be found at the same thresholds: those where $G(\tau)=F(\tau)$. In particular our result about the dynamics (Theorem 4) immediately applies with the appropriate sufficient condition, and the Gaussian case has the same equilibra with the same stability.
Beyond consistent threshold strategies.
The above analysis is restricted to our weaker notion of equilibrium in the space of consistent threshold strategies. It is unclear whether these remain equilibria if agents can use any consistent strategy, or if agents can use any four threshold strategies. Even more broadly, one could consider joint-task strategies $\sigma:\mathbb{R}^{4}\to\{H,L\}^{4}$, which map all four signals simultaneously to all four reports. While not shown in Kong [2020, 2024], in the binary report setting the DMI mechanism turns out to satisfy the even stronger property that truthfulness is an equilibrium among all joint-task strategies. We prove this claim in Theorem 9 in § C. We then show that truthfulness is not an equilibrium for real-valued reports: in many cases the best response flips one or two of the truthful reports to increase the chance of a nonzero payoff in $M_{\textrm{DMI}}$.
5 Robust Bayesian Truth Serum
We now study the Robust Bayesian Truth Serum (RBTS) mechanism proposed by Witkowski and Parkes [2012]. RBTS is inspired by the Bayesian Truth Serum (BTS) [Prelec, 2004]. However, while incentive alignment under BTS requires an arbitrarily large number of agents, RBTS guarantees truthfulness for $n\geq 3$ agents.
RBTS is a non-minimal mechanism: each agent $i$ submits both an information report $r_{i}\in\{L,H\}$ and also a prediction report $p_{i}\in[0,1]$. The prediction report represents agent $i$’s belief about the frequency of $H$ reports from all agents. The mechanism incentivizes prediction reports using the Brier score $S(p,r)=2p\mathbf{1}[r=H]-p^{2}$ (also known as the quadratic score), which is an example of a strictly proper scoring rule [Brier, 1950, Gneiting and Raftery, 2007]. RBTS randomly picks a reference agent $j$ and peer agent $k$, and pays agent $i$
where $t(L)=-1$ and $t(H)=1$, and $\delta>0$ is chosen by the mechanism.
We refer to the first term in Equation 17 as the information score, and the second as the prediction score. Under binary signals and reports, truthfulness over both the information and prediction report forms a strict Bayes-Nash equilibrium.
5.1 Equilibrium Characterization
Proposition 6.
$\tau^{*}=\pm\infty$ are threshold equilibria under RBTS.
Proof.
We also observe necessary and sufficient conditions for non-infinite equilibria $\tau$, relative to the function $G$. Let $Q(x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(x;x^{\prime})$, and let $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ be the expected conditional probability of another agent reporting $L$ over signal $x$.
Theorem 6.
Let a finite threshold $\tau$ be given and $P(\tau;x)$ be continuous over $x$; also assume $\beta(x)$ is continuous over $x$. If $\tau\in\mathbb{R}$ is threshold equilibrium under the RBTS mechanism then $Q(\tau)=G(\tau)$. Conversely, if $Q(\tau)=G(\tau)$ and $P(\tau;x)-\bar{P}(\tau;x)$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the RBTS mechanism.
Proof.
For necessity, first note that since $\beta(x)$ and $P(\tau;x)$ are continuous over $x$, $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ is also continuous over $x$ by an application of the Dominated Convergence Theorem. Let $f(\tau;x)=P(\tau;x)-\bar{P}(\tau;x)$. Since $f(\tau;x)$ is continuous, we must have $\lim_{x\to\tau^{+}}f(\tau;x)=\lim_{x\to\tau^{-}}f(\tau;x)=f(\tau;\tau).$ Further, $\lim_{x\to\tau^{-}}f(\tau;x)\geq 0$ and $\lim_{x\to\tau^{+}}f(\tau;x)\leq 0$ by Conditions 18 and 19, so $f(\tau;\tau)=0=G(\tau)-Q(\tau)$.
While the exact crossing condition has changed from $G(\tau)=F(\tau)$ to $G(\tau)=Q(\tau)$, Theorem 6 paints a similar picture to what we saw for DG. Best responses arew now more complex to compute, so we do not treat the general dynamics formally, but we can provide a comparison of their behavior in the Gaussian case.
5.2 Gaussian Model
We return to our Normal example, where each agent receives a noisy version of a Gaussian signal. Leaving the technical details to § D, we first note that the same equilibria occur as the previous mechanisms.
Corollary 5.

### v2

@@ 3.3 Determinant-based Mutual Information (DMI) Mechanism
In § D, under some constraints on the form of strategies over each of the 4 task signals, we show that DMI has the same necessary condition for equilibrium as DG. So, modulo a slightly different sufficient condition, its equilibria can be found at the same thresholds: those where $G(\tau)=F(\tau)$. In particular, our result about the dynamics (Theorem 5) immediately applies with the appropriate sufficient condition, and the Gaussian case has the same equilibra with the same stability.
4 Distributions beyond the Gaussian
We now aim to numerically explore existence and dynamics of equilibria with real-valued signals across mechanisms in settings other than the noisy Gaussian model. We consider two natural departures from the “nice” properties of Gaussians, skewness and multimodality.
4.1 Skewness
We extend our observations to asymmetric signal distributions by adapting the same noisy signal model as our running Gaussian example, but with a base distribution which is skewed. Formally, each agent $i$ receives a signal $X_{i}$ such that $X_{i}=Z+Z_{i}$, where $Z$ is the skewed Normal distribution with mean 0 and variance 1, and $_{i}\sim N(0,1). ZZ$ has another parameter, $\alpha$, which controls the skewness of the distribution. The magnitude of $\alpha$ increases the magnitude of skewness; moreover, $\alpha<0$ leads to a left-skewed distribution, and $\alpha>0$ a right-skewed distribution. We study how threshold equilibria move as a function of the parameter $\alpha$.
As depicted in Figure 5, we find a unique finite equilibrium under OA, DG (and DMI, by our theoretical results), and RBTS which shifts toward the longer tail, i.e. the thresholds move from negative to positive as the skewness moves from left to right. We recover the equilibrium at $\tau=0$ in the original Gaussian model when $\alpha=0$. Thus we expect when there is asymmetry in the distribution of signals, agents in DG, DMI, and RBTS will settle at a threshold that balances out the asymmetry in signal mass. While we observe the same high-level dynamics for OA, these equilibria are unstable and so we still expect agents will move toward an uninformative consensus.
4.2 Multimodality
Next we explore the effects that a distribution with multiple peaks has on the equilibria that occur under our signal model. Consider a classic mixed Gaussian setting where each agent’s signal $X_{i}$ is received i.i.d. from one of $K$ Gaussian components, each with mean $\mu_{k}$ and variance $\sigma_{k}^{2}$. There is an underlying parameter $z\sim\text{Categorical}(\pi)$ for $\pi\in\Delta^{K}$ that determines which Gaussian component all the signals come from. For the purposes of our experiments, we fix $\pi$ to be the uniform distribution over the $K$ components. We numerically calculate the functions $F$, $G$, and $Q$, and check for the sufficient and necessary conditions of our theorems, to find equilibria and their stability. We set the precision of each component to be the same, and then vary this precision parameter to generate bifurcation graphs for $K=2$ and $3$ (see § E for $K=4$).
Figure 4: $F$, $G$, and $Q$ over signals $x$ in the two-component Gaussian mixture setting with means 1,1-1,1 and precision 2.
Figure 5: Bifurcation diagrams for two- and three-component Gaussian mixtures. (Left) Two-component case, with means 1,1-1,1: OA (solid) exhibits three equilibria at high precision, while DG (dashed) maintains a single stable equilibrium at $\tau=0$. (Right) Three-component case with means 2,0,2-2,0,2: at high precision, OA (solid) has five equilibria, with two stable at $\tau=\pm 1$; DG (dashed) has three equilibria, with stability reversed from the two-component OA case.
We first observe that the number of equilibria increases by two each time we add a Gaussian component (Figure 5) for both OA and DG. OA begins with three equilibria for large enough precision under the two-component model, with $\tau=0$ emerging as a nontrivial stable point. In general, for high enough separation between the two modes, stable equilibria exist beyond trivial agreement, and the area of starting conditions that lead to these equilibria increases. One can view this phenomenon as a reflection of the original binary signal model: with enough separation between the two modes of the richer signal space, we essentially return to a setting where Output Agreement is useful because the “signal” is effectively which mode has been chosen. The pattern continues more generally for $K$ components, with stable threshold equilibria emerging between pairs of modes for high enough precision values.
For DG, $\tau=0$ remains the unique stable equilibrium for two components, while two stable equilibria emerge as the precision increases at $\pm 1$, between the distribution’s modes, for three components. It follows that if signal noise is sufficiently small and the underlying distribution is multimodal (with at least three modes), the designer has more choice in how agents map information to reports. E.g., in peer grading, if the quality distribution has three modes, the designer can choose between thresholds separating low and high marks at the first or second mode. On the other hand, the designer does not have control or knowledge of the underlying signal distribution.
RBTS and DMI.
We view equilibria in the snapshot of a specific precision value in Figure 5. The intersection of $G$ with 1/2 corresponds to equilibria in OA, with $F$ corresponds to equilibria in DG (and DMI), and with $Q$ corresponds to equilibria in RBTS. The function $Q$ closely tracks $G$ for high enough precision values. We therefore find that the same stability and number of equilibria occur in RBTS as in DG. However, we note that bifurcation (the increase in equilibria) occurs at lower precision values for RBTS (see § E for visualization). From a design perspective, a lower bifurcation point in RBTS is nice because there is a wider set of signal distributions allowing for some choice over stable thresholds.
5 Flexibility of a Richer Report Space
While a single, inflexible threshold equilibrium at the median exists under the Gaussian model, we found in Section 4 that multiple stable threshold equilibria emerge when the underlying signal distribution is multimodal with low individual signal noise. However, in peer prediction settings the designer cannot control, and often cannot diagnose, modality of the signal distribution. Moreover, when the distribution is unimodal, our results remain negative. A natural next question is whether these negative results persist when the report space is larger: i.e., what if the designer aims to employ a peer prediction mechanism designed for three reports instead of two?
In this section, we study an extension of DG to non-binary reports, Correlated Agreement (CA) [Shnayder et al., 2016a]. Under the same model of equilibria and dynamics, we find issues of inflexibility persist for any parameters of the mechanism: if there are $m$ possible reports under the Gaussian model, there only exists a single nontrivial, stable threshold equilibrium $\tau\in\mathbb{R}^{m-1}$. While the outlook is still negative in this case, we uncover a new peer prediction framework with more flexibility by leveraging our results for the non-binary report setting. Specifically, if the designer intends for agents to map their signals to a binary report, what if we first ask for more? E.g., the designer uses CA with a report space $\{1,2,3,4\}$ of size $m=4$, and is then able to generate a mapping back to $\{L,H\}$. If there exists a three-dimensional stable threshold equilibrium at, say, (0.5,0,0.5)(-0.5,0,0.5), the designer can choose between a binary threshold of 0.5-0.5, 0, or $0.5$ through this mapping. We formalize this report-mapping framework in § 5.4 and discuss how the designer can set CA parameters in order to guarantee these choices exist over thresholds.
5.1 Correlated Agreement
We begin by studying the non-binary report setting under a natural extension of DG called Correlated Agreement [Shnayder et al., 2016a]. At a high level, CA gives the designer flexibility in allowing agents to receive payments when their peers submit reports that differ from their own.
Let $m]=\{1,2,\dots,m\}$, and let the report space be $\mathcal{R}=[m]$. Each agent $i$ still calculates a report $r_{i}\in\mathcal{R}$ according to deterministic strategy $\sigma_{i}:\mathbb{R}\to\mathcal{R}$. Now we consider a threshold vector $\tau\in\mathbb{R}^{m-1}$ with $\tau_{1}<\tau_{2}\dots<\tau_{m}$, where a threshold strategy is defined as follows:
$\sigma^{\tau}(x)=\begin{cases}1,&x\leq\tau_{1}\\ 2,&\tau_{1}<x\leq\tau_{2}\\ \ldots\\ k,&\tau_{k-1}<x\leq\tau_{k}\\ \ldots\\ m,&x>\tau_{m-1}.\end{cases}$
For convenience we define $\tau_{0}=$-$\infty$ and $\tau_{m}=\infty$.
In CA, the designer chooses a symmetric, binary $m\times m$ score matrix $\Delta$. While Shnayder et al. [2016a] suggest initializing $\Delta$ as the sign matrix of correlations between signal bins, this underlying structure is often unknown, and furthermore depends on the current thresholds used by agents under our dynamics model. Thus we allow the designer to fix the score matrix beforehand. Then if agents $i$ and $j$ submit reports $r_{i}$ and $r_{j}$ for some shared task, and a peer reports $r_{k}$ on a distinct, i.i.d. task, agent $i$ receives payment $\Delta(r_{i},r_{j})-\Delta(r_{i},r_{k})$.
As in DG, if all agents are using a symmetric threshold strategy $\sigma^{\tau}$, agent $i$’s expected penalty for reporting $\ell$ is $\pi_{\ell}$, where $\pi_{\ell}=\sum_{k\in[m]}\Pr[\tau_{k-1}<X\leq\tau_{k}]\Delta(\ell,k)$. That is, $\pi_{\ell}$ is the expected score over the prior of report $\ell$ against another agent using the symmetric strategy on a randomly selected task. Thus we work with the following formulation of CA with the same expected payment:
$M_{CA}(r_{i},r_{j},\pi)=\Delta(r_{i},r_{j})-\pi_{r_{i}}.$ (11)
Note that when using the identity matrix as $\Delta$, CA exactly corresponds to extending the DG mechanism to $m$ reports (and in the case of binary reports, the identity matrix is the only valid score matrix, meaning CA is equivalent to DG). However, for other matrices $\Delta$, an agent’s expected utility now depends on the probability of any report $j$ such that $\Delta[i,j]=1$.
Under more general monotonicity conditions over the utility functions of each report $k$ (see Condition 3 in § F), we can derive a similar characterization of equilibria by equating utilities of adjacent reports at each threshold. As their forms are intuitive extensions of our analysis for binary-report settings, we leave the formal equilibrium characterization to § F. We also note that it is possible for thresholds to collapse, e.g. $\tau_{1}=\tau_{2}=0$. In this case the score matrix also collapses to a smaller report space, and we can use our characterization to check for equilibria with our new effective number of thresholds.
Dynamics.
If our monotonicity condition (Condition 3) hold for a fixed threshold $\tau$, it immediately follows that the best response to a threshold strategy $\tau\in\mathbb{R}^{m-1}$ is also a threshold $\hat{\tau}\in\mathbb{R}^{m-1}$. We can therefore study the same dynamics model, with $\dot{\tau}=\hat{\tau}-\tau$. As $m$ increases, analytically studying the dynamics of threshold strategies becomes more unwieldly, so in this section we numerically simulate the system for different initial conditions and visualize the vector field to understand which equilibria are stable or unstable. We focus on the Gaussian model, normalizing the base signal so its variance is $a^{2}=1$. After numerically guaranteeing that Condition 3 holds in each case, we vary the individual signal precision ($1/b^{2}$) and study how equilibria change.
Figure 6: Vector fields of thresholds $\tau\in\mathbb{R}^{2}$ for CA with score matrix $\Delta_{I}$, $m=3$, and precision 11 (left) and 44 (right). The basins of attraction for each stable equilibrium are indicated by color.
Figure 7: Nontrivial stable threshold equilibria for CA with score matrix $\Delta_{I}$ (solid) or $\Delta_{T}$ (dotted) for $m=3$ (left) and $m=4$ (right), varying precision of signals $Z_{i}$ under the Gaussian model.
5.2 Diagonal score matrix (DG)
Let $\Delta=\bm{I}$ be the identity matrix, so that CA serves as an extension of DG to non-binary reports. For any precision value, when $m=3$ we find there exists a stable equilibrium at a point (c,c)(-c,c) for some $c\geq 0$. Figure 7 shows how the two thresholds at c-c and $c$ change with signal precision. At small precision values, $c$ collapses to 0, so that we return to the same equilibrium as the binary report case (e.g., report ‘1’ for all signals below 0, and ‘3’ otherwise). Because the posterior is flatter with longer tails, there is enough mass above any threshold $\tau_{2}>0$ relative to the prior that agents with signals $\tau_{2}-\epsilon$ will prefer reporting ‘3’, and vice versa for $\tau_{1}<0$ with signals $\tau_{1}+\epsilon$. However, for larger precision values, shorter tails allow the value of $c$ to move away from zero.
Figure 7 illustrates that alongside this nontrivial stable equilibrium (blue), there also exist stable equilibria at $(-\infty,0)$ and $(0,\infty)$. Again, these equilibria effectively collapse thresholds to 0, e.g. agents report ‘2’ when receiving signals below 0 and ‘3’ otherwise in the former case. These collapsed, stable equilibria are separated from the center point (c,c)(-c,c) by two hyperbolic equilibria (red). Intuitively, the basin of attraction for such collapsed equilibria correspond to settings with enough positive correlation between two adjacent signals far enough in the tails that agents are incentivized to collapse their meanings together. In general, then, we observe that there still only exists (at most) one nontrivial stable equilibrium beyond a collapsed threshold of 0, and the designer has no choice in where this point is.
The pattern continues when $m=4$ in Figure 7 (indicated by the solid lines), with higher correlation coefficients leading to nontrivial stable threshold equilibria at some symmetric $\tau=(-c,0,c)$. Again, there exist several stable equilibria collapsing the threshold to 0, now at $(-\infty,0,0)$, $(0,0,\infty)$, $(-\infty,-\infty,0)$, and $(0,\infty,\infty)$, separated by saddle points from the center equilibrium at $\tau$. Our observations generalize for larger $m$, with at most one nontrivial stable equilibrium of dimension m1m-1 emerging for high enough precision (see § F.3 for more details).
5.3 Tridiagonal matrix
We also study equilibria and their stability under different precison values when the score matrix corresponds to the tridiagonal matrix $\Delta_{D}$, where $\Delta_{T}[i,j]=1$ when $i$ and $j$ are adjacent. E.g., for $m=3$ and $m=4$ respectively, $\Delta_{T}$ looks like
$\begin{pmatrix}1&1&0\\ 1&1&1\\ 0&1&1\end{pmatrix}\qquad\text{and}\qquad\begin{pmatrix}1&1&0&0\\ 1&1&1&0\\ 0&1&1&1\\ 0&0&1&1\end{pmatrix}$.\\
When signal precision is lower, it is reasonable to expect this mechanism is robust against thresholds collapsing because agents receive payments for a neighborhood of signals which may be more likely under their posteriors. We begin by noting that for a report space of size $m=3$, the expected reward under the tridiagonal matrix for report $2$ is zero, and agents are always incentivized to collapse this middle signal and return to the binary report case (see § F.2 for a formal analysis). [^5] This perhaps reflects the fact that in the original CA setup, the score matrix cannot be tridiagonal when $m=3$ because a signal cannot be positively correlated with all signals (as the second row indicates).
We visualize the setting of $m=4$ with the score matrix $\Delta_{T}$ in Figure 7. Unlike the diagonal score matrix, here there exists a unique, non-zero, stable equilibrium $\tau\in\mathbb{R}^{3}$ even when the precision is low (indicated by the dotted lines). For example, when the precision is $1$, a stable equilibrium emerges under $\Delta_{T}$ with $\tau\approx(-0.73,0,0.73)$. While a collapse to 0 does not occur, agents will still always converge to this unique set of thresholds, and thus the designer still has no flexibility in choice over equilibria. We numerically confirm this behavior for higher $m$ in § F.3.
5.4 Report mapping framework
The previous subsections show that with non-binary report spaces, employing CA with either score matrix leads to the same inflexibility of equilibria. So is there any hope in developing a mechanism that provides the designer more choices in how agents map signals to reports? To make progress toward this goal, we propose a framework which leverages our results for non-binary reports.
@@ 5.4 Report mapping framework
(Report mapping framework) We define the report mapping framework as follows.
  1. 1.
    Run the Correlated Agreement mechanism over $m$ raw reports with fixed score matrix $\Delta$. E.g., if $m=3$, agents submit one of three reports in $\{1,2,3\}$.
  2. 2.
    Map each agent’s raw report to a final binary report of $L$ or $H$.
Intuitively, given our results in the previous section, running step 1 with $m$ reports will lead to a threshold equilibrium of dimension m1m-1 in some regimes. For example, when $m=3$ and the precision is 2, there is a unique stable threshold equilibrium $\tau\in\mathbb{R}^{2}$ when $\Delta=\bm{I}$ with $\tau\approx(-0.5,0.5)$, so that agents report ‘1’ if their signal is below 0.5-0.5, ‘2’ if their signal is between 0.5-0.5 and $0.5$, and ‘3’ otherwise. In step 2, the designer can then choose from either 0.5-0.5 or $0.5$ as a final threshold: e.g., a threshold of 0.5-0.5 is simulated by mapping reports of ‘2’ and ‘3’ to $H$ and reports of ‘1’ to $L$. This is in contrast to our results using binary-report peer prediction mechanisms directly, where agents inevitably converge to a threshold of 0.
Using the identity matrix $\Delta_{I}$.
As previously mentioned and shown in Figure 7, when the precision is low enough equilibria still collapse to a single threshold at 0 under score matrix $\Delta_{I}$, meaning there is no inherent benefit in running the report mapping framework for $m=3$ or $m=4$. However, when precision is high, the designer can choose from three different thresholds. For example, if $m=3$ and the precision is $\geq 3/2$, a nontrivial threshold equilibrium emerges of the form (c,c)(-c,c), limiting to around (0.5,0.5)(-0.5,0.5); and for $m=4$, if the precision is $\gtrapprox 4$, a nontrivial threshold equilibrium emerges of the form (c,0,c)(-c,0,c) limiting to around (0.75,0,0.75)(-0.75,0,0.75). As evident in Figure 7, the basin of attraction for these nontrivial equilibria also increases in size as precision grows, so that there is a larger margin of error for the designer in estimating the initial threshold.
We note that the bifurcation point at which equilibria no longer collapse to 0 is higher for $m=4$ than for $m=3$, intuitively because the amount of probability per report bin under the signal distribution decreases and exacerbates agents near thresholds reporting against them. We numerically verify that this pattern continues, with the minimum precision required for a non-zero, stable equilibrium $\tau\in\mathbb{R}^{m-1}$ to emerge increasing over $m$; see § F.3 for visualization. Thus there is a tradeoff when using $\Delta_{I}$: while larger values of $m$ increase the number of potential thresholds in the report-mapping framework that the designer can choose from, the signal precision required for a non-zero, stable equilibrium $\tau\in\mathbb{R}^{m-1}$ to emerge in the first place is also increasing over $m$.
Using the tridiagonal matrix $\Delta_{T}$.
As a reminder, when $m=3$ our only nontrivial scoring matrix option is the identity. But for $m=4$, employing the report-mapping framework when using the score matrix $\Delta_{T}$ is useful, even for low precision values. Specifically, there exists a nontrivial, unique threshold equilibrium even when agent information is highly noisy. Intuitively, the correlation between an agent’s signal and adjacent bins is high enough under $\tau$ that, even under higher signal noise, they are incentivized to report truthfully according to the bin their signal is in. In these regimes, then, the mechanism designer is able to choose between three different threshold mappings of signals to final reports (c-c, 0, or $c$ for some $c>0$).
This pattern continues for larger $m$, with a single stable threshold equilibrium $\tau\in\mathbb{R}^{m-1}$ existing for any precision level. We note that there is a tradeoff for larger report spaces: as $m$ grows and precision limits to 0, the outer thresholds move toward $\pm\infty$. However, this transition to a smaller-dimension equilibrium occurs at a significantly lower precision level than the equilibrium collapse under the diagonal matrix $\Delta_{I}$. We direct the reader to § F.3 for a more detailed discussion.
Takeaways.
In practice, then, using the report-mapping framework with the tridiagonal matrix $\Delta_{T}$ is useful when the designer is either unsure of underlying agent signal precision, or believes the precision is low. For example, in peer grading the designer may believe there is high variability in how individual graders read an essay due to skill or lack of effort, or there is high subjectivity or disagreement about the quality of a submission. When the designer has confidence that signal precision is higher, e.g. graders evaluate essay qualities more consistently, the diagonal matrix $\Delta_{I}$ can also be used to ensure there exist several threshold options for the designer.
We additionally note that while this section focuses on settings where the principal is seeking out binary reports, based on our results the report-mapping framework could be adapted to any final report space, as long as the raw report space is set to be larger.
6 Discussion
We initiate the study of peer prediction when signals are finer-grained than reports. To start, we characterize behavior in several binary-report peer prediction mechanisms when signals are real-valued. While all these mechanisms have various truthfulness guarantees in the binary signal setting, in our continuous signal model we find that the notion of truthfulness breaks. Specifically, when agents map signals to reports according to a real-valued threshold, under dynamics we find equilibria rarely stay at the initial threshold a mechanism designer sets. Instead, agents will often stabilize at a different (and potentially uninformative) threshold according to the specific payment scheme. We demonstrate that this inflexibility of equilibria extends to other general settings where signals are finer-grained than reports, e.g. when the report space is non-binary or when the signal space is finite (see below).
Motivated by these negative results, we present a new report-mapping framework which allows the principal more flexibility in setting threshold equilibria. By using the Correlated Agreement mechanism to elicit raw reports from a larger space, the designer has the flexibility to choose several different mappings between these reports and $\{L,H\}$ depending on the norms with which they prefer agents to interpret their signals.
Our results imply that the standard modeling assumptions in peer prediction literature miss important nuances, and these holes can propagate to real changes in how we expect agents to behave in practice. In particular, models of peer prediction have often taken as given an arbitrary joint distribution $P$ over pairs of signals drawn from a finite set. This suggests that we should be able to choose an arbitrary meaning for each signal and then classic peer prediction should “just work” for the resulting distribution. In contrast, our results show that classic peer prediction is substantially more inflexible as currently used in practice. Fortunately, our report-mapping framework avoids this collapse of information to a single equilibrium, by asking agents to map their information to a finer list of reports first. We discuss several extensions of our results and avenues for future work below.
Beyond real-valued signals.